Find four consecutive integers such that twice the sum of the two greater integers exceeds three times the first by 91.

Thank you so much!!!!!
Kelli

Answer by venugopalramana(3286)

Find four consecutive integers such that twice the sum of the two greater integers exceeds three times the first by 91.
LET N,N+1,N+2,N+3 BE THE 4 INTEGERS
SUM OF 2 GREATER INTEGERS = N+2+N+3=2N+5
TWICE THE SUM = 2(2N+5)=4N+10
3 TIMES THE FIRST = 3N
4N+10=3N+91
N=81
SO THE 4 INTEGERS ARE 81,82,83,84